Novus can raise equity capital in its domestic equity market, or in the global equity markets.  Novus's equity beta in the domestic market is 0.95.  In the global equity markets, which have a higher expected return, Novus has a lower beta of 0.625. The risk-free rate in either the domestic or global securities market is 3.5%, while the domestic market has an expected market return of 10% and the global market has an expected return of 12.5%.

a)      Calculate Novus's cost of Equity for both portfolio data sets

Component (Symbol)

Domestic Portfolio

Global Portfolio

Risk-free Rate (krf)

3.5%

3.5%

Market return (km)

10.0%

12.5%

Beta (β)

0.950

0.625

Cost of equity (ke)

b)   If Novus's debt/total capitalization ratio increases from 35% to 45%, given its cost of debt of 4% and its tax rate of 20%, calculate the WACC for each portfolio set before and after the change in capital structure.

Portfolio Set

35% Debt WACC

45% Debt WACC

Domestic

Global
Answer #1

Answer A:

Cost of Equity = Risk-free rate + Beta (Market return - riskfree rate)

--------------------

Cost of Equity in Domestic market :

Risk free in domestic market : 3.5% , market return = 10% , beta = 0.950

Cost of Equity = 3.5 + 0.95( 10-3.5) = 9.675%

-------------

Cost of Equity in global market :

Risk free rate = 3.5%, market return = 12.5 , beta = 0.625

Cost of Equity = 3.5 + 0.625(12.5 - 3.5)

= 9.125%

----------------------------

Answer B

WACC = WeRe + WdRd(1-t)

where We = weight of equity , Wd = weight of debt

Re = cost of equity , Rd = cost of debt , t = tax

WACC in domestic portfolio,  when debt = 35% ,

Wd = 0.35 , We = 1-0.35 = 0.65

Cost of debt = 4% , t = 20% , cost of equity in domestic portfolio = 9.675%

WACC = 0.65*0.09675 + 0.35*0.04*(1-0.2)

= 7.409% = 7.41% (rounding off)

WACC in global portfolio, when debt = 35%,

cost of equity in global portfolio = 9.125% , rest data remains same.

WACC = 0.65*0.09125 + 0.35*0.04*(1-0.2)

= 7.051% = 7.05%(rounding off)

----------------------------------

WACC in domestic portfolio,  when debt = 45% ,

Wd = 0.45 , We = 1-0.45 = 0.55

Cost of debt = 4% , t = 20% , cost of equity in domestic portfolio = 9.675%

WACC = 0.55*0.09675 + 0.45*0.04*(1-0.2)

=6.761% = 6.76%(rounding off)

-----------

WACC in global portfolio, when debt = 45%,

cost of equity in global portfolio = 9.125% , rest data remains same.

WACC = 0.55*0.09125 + 0.45*0.04*(1-0.2)

= 6.459% = 6.46% ( rounding off)

Hope this helps you!!!!